OpenStudy (anonymous):

Write the equation of the line which passes through (5, –2) and is parallel to x = 4. (1 point)

OpenStudy (nikita2):

x=5

OpenStudy (anonymous):

finding -2=1(5)+c c=-7 y=-2(4)+-7 y=-15

OpenStudy (nikita2):

But line y=-15 does not pass through (5,-2)

OpenStudy (anonymous):